A car traveled at 60 miles per hour for 2½ hours. If one inch equals 20 miles on a
map, how far has the car traveled on the map?

Found 2 solutions by corncob, mananth:
Answer by corncob(8)

To start you will multiply 60*2.5 to get how many miles the car has traveled.
(The 60 is the miles per hour and the 2.5 is the hours it traveled.)
You should end up with 150 miles.
Then you will divide 150 by 20 (The 20 represents the miles per inch)
150/20=7.5
So, the answer is 7.5 inches

Answer by mananth(16946)

60 mph for 2 1/2 hours
2 1/2= 2.5 hours
..
distance = r*t
d=60 * 2.5
d=150 miles.
..
MAP
20 miles = 1 "
150 miles = 150/20 = 7.5 inches
...
m.ananth@hotmail.ca